Find an orthogonal matrix P whose first row is u1=(1/3, 2/3, 2/3)
1
Expert's answer
2020-12-15T03:14:07-0500

Since we need matrix A є M3x3 (F) which is orthogonal, we want orthonormal basis for F3 whose first vector is given row "(1\/3; 2\/3;2\/3)" (because row of an orthogonal matrix M form an orthonormal basis for Fn).

Let "\gamma = {\{(1/3, 2/3, 2/3), (1,0,0), (0,1,0)}\}" and it be ordered basis for F3 with given row as first element.

Set "\omega1 = (1/3,2/3,2/3), \omega2 = (1,0,0), \omega3 =(0,1,0)" and use Gram-Schmidt orthogonalization process to get orthogonal basis for F3 .


"\upsilon1 = \omega1 = (1/3, 2/3, 2/3)"

"\upsilon2 = \omega2 - \upsilon1 * <\omega2, \upsilon1>/<\upsilon1, \upsilon1> ="

"= (1,0,0) - (1\/3,2\/3,2\/3)*(1\/3)\/(1\/9+4\/9+4\/9)"

"=(1-1\/9,0-2\/9, 0-2\/9) = (8\/9, -2\/9, -2\/9)" "= (1,0,0 ) - 1\/3*(1\/3,2\/3,2\/3)"

"\upsilon3 = \omega3 - \upsilon1 * <\omega3, \upsilon1>/<\upsilon1, \upsilon1> -\upsilon2 * <\omega3, \upsilon2>/<\upsilon2, \upsilon2> ="

"= (0,1,0) - (1\/3, 2\/3, 2\/3)*(2\/3)\/(1\/9+4\/9+2\/9) - (8\/9,-2\/9,-2\/9)*(-2\/9)\/(64\/81+4\/81+4\/81)""=(0,1,0) - 2\/3 *(1\/3,2\/3,2\/3) +1\/4*(8\/9, -2\/9,-2\/9)" "="

"= (0 - 2\/9+8\/36, 1-4\/9-2\/36, 0 - 4\/9-2\/36)"

"= (0, 1\/2, -1\/2)"

"\upsilon1,\upsilon2, \upsilon3" is now orthogonal basis which only needs to be normalized.

Define "ui = \upsilon i/||\upsilon i|| for" i=1,2,3

"\upsilon1" is normal yet, so u1 = "\upsilon1" = (1/3, 2/3, 2/3)

"u2 = \upsilon2/||\upsilon2||"

"=(8/9,-2/9,-2/9)/\sqrt{\smash[b]{8/9*8/9+(-2)/9*(-2)/9 +(-2)/9*(-2)/9}}"

"= (8/9,-2/9,-2/9)/(2*\sqrt{\smash[b]2}/3)"

"= (2*\sqrt{\smash[b]2}/3, -\sqrt{\smash[b]2}/6, -\sqrt{\smash[b]2}/6)"

"u3 = \upsilon3/||\upsilon3||"

"=(0,1/2,-1/2)/\sqrt{\smash[b]{0+1/2*1/2 +(-1)/2*(-1)/2}}"

"= (0,1/2,-1/2)/(1/\sqrt{\smash[b]2})"

"= (0, \sqrt{\smash[b]2}/2, -\sqrt{\smash[b]2}/2)"

Now we have orthonormal basis

"\{ (1/3, 2/3, 2/3), (2*\sqrt{\smash[b]2}/3, -\sqrt{\smash[b]2}/6, -\sqrt{\smash[b]2}/6) ,(0, \sqrt{\smash[b]2}/2, -\sqrt{\smash[b]2}/2) \}"

and the only one thing we need to do is to define the matrix with rows being the vectors from that basis:

"\begin{pmatrix}
 {1 \over 3} & {2 \over 3} & {2 \over 3} \\
 {2* \sqrt{\smash[b]2}\over 3} & {-\sqrt{\smash[b]2} \over 6} & {-\sqrt{\smash[b]2} \over 6}\\
 0 & {\sqrt{\smash[b]2} \over 2} & {-\sqrt{\smash[b]2} \over 2}
\end{pmatrix}"